OpenStudy (anonymous):

In a game with 2 dice, you win if you get a sum of 7 or 11 or roll doubles. a. what are the chances against winning this game? b. explain how you solved this question.

OpenStudy (mathmate):

There are 6 ways to get a 7, and 2 ways to get an 11, all out of 36 outcomes. Chances of winning game is therefore (7+2)/36=1/4, and chances against winning is 1-1/4=3/4 (or odds 3:1) Either count by enumeration, or by drawing a 6x6 table and counting the outcomes.